Operation Summary for February 12, 2010 - February 15, 2010
Friday Feb. 12
The day shift began with Operations monitoring the Tevatron (TeV) store 7603 (119.04E30) and the Recycler stash (5.8E10), with the Antiproton Source stacking (31.0E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beam.
During the day shift, at 3:30 PM, store 7603 quenched at sector A3. Operations prepared the TeV A-E sectors and Transfer Hall for access so that Mechanical Support personnel could hang portable vacuum pumps ( turbocarts) in the area. The TeV quench recovery continued throughout the shift. Operations reported that RF technicians changed out a MI RF station (MIRF9) modulator.
During the evening shift, the A3 quench recovery continued. Vacuum techs accessed the TeV and Transfer Hall areas again to install more turbocarts. Operations had current on the TeV bus by 11:46 PM.
Saturday Feb. 13
The midnight shift began with the TeV turn-on beginning, with Operations monitoring the stash (318.3E10), with Pbar stacking (39.9E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations put the TeV into a dry squeeze for the rest of the shift. Operations reported problems with the following Linac devices that kept beam off throughout the shift: LRF2 IPA, the Klystron buncher, and the Klystron Debuncher LCW water-skid, which required more support personnel for assistance. There was only about 30 minutes of beam to users during this shift.
During the day shift, Linac repairs continued while the TeV sat in a dry squeeze. The MINOS absorber and alcoves, along with M01, were accessed. Linac personnel completed their repairs at 1:30 PM. Operations reported that stacking was interrupted by many Pbar Debuncher kicker trips; this problem was traced to a high voltage power supply, which was replaced. Operations put the TeV into shot setup at 3:50 PM and established store 7606 at 5:14 PM with an Average Initial Luminosity (AIL) of 248.8E30.
During the evening shift, Operations established beam to MTest at 5:48 PM and terminated the beam at 6:07 PM.
Sunday Feb. 14
The midnight shift began with Operations monitoring store 7606 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to MTest at 4:34 AM. Operations reported that store 7606 aborted at 5:04 AM when the hardware (interlock) loop interface chassis at TeV sector A2 failed and needed to be changed out. Operations had the TeV turn-on underway by the end of the shift. Operations reported problems in the Booster East Gallery due to a loss of TCLK signal that caused about an hour of downtime. The problem went away during troubleshooting. Operations reported a TeV data acquisition (DUE01) hardware problem; hardware techs spent time diagnosing the problem.
During the day shift, Operations put the TeV into shot setup at 10:47 AM and established store 7610 at 12:14 PM with an AIL of 296.6E30. Operations reported several NUMI beam permit trips due to a power supply readback (V100E) drifting out of tolerance. Jasmin experimenters accessed M01 twice during the shift.
During the evening shift, Operations disabled the MTest beam at 6:06 PM.
Monday Feb. 15
The midnight shift began with Operations monitoring store 7610 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 7610 at 3:13 AM and put the TeV into shot setup. Operations established store 7611 at 4:36 AM with an AIL of 256.4E30. Operations established beam to MTest at 5:08 AM. A Booster RF station (BRF9) tripped off on a LCW flow switch indication; the station’s filaments needed to time-in once the station was reset.
The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to repair TCLK. MTA (MuCooL) will take some beam this afternoon.
Comments and Suggestions
|last modified 02/19/2010 email Fermilab|
|Security, Privacy, Legal|